6.movie tickets for an adult and three children cost $20. An adults ticket costs $2 more than a child's ticket. find the cost of an adult's ticket.

Answer by evhil_mahabahkah(3)

Let x= cost of a child's ticket
x+2= cost of a adult's ticket
3x= cost of 3 child's tickets
3x+x+2 = $20
4x+2 = $20
4x = $20-$2
4x/4 = $18/4
x = $4.50
therefore : x+2= $4.50 + $2 = $6.50